OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.

OpenStudy (lgbasallote):

use point slope formula \[y-y_1 = m(x-x_1)\] does that help?

OpenStudy (lgbasallote):

or do you need more help?

OpenStudy (anonymous):

yes please

OpenStudy (zzr0ck3r):

y = mx+b we know m y = 2x+b now plug in the values -2 = 2(3) + b -2 = 6+b -2-6 = b b = -8 so now what is the equation?

OpenStudy (lgbasallote):

in my equation..you just plug in directly what you have m = slope = 2 x1 is the x coordinate = 3 y1 is the y coordinate = -2 so plug it in \[y - y_1 = m(x-x_1)\] \[y-(-2) = 2(x-4)\] now simplify that equation to get your equation of the line

OpenStudy (zzr0ck3r):

im pretty sure @lgbasallote had a typo and meant y-(-2) = 2(x-3)... just to let you know.

OpenStudy (lgbasallote):

uhh yeah thanks for that...didnt notice it